OpenStudy (anonymous):

Two angles are complementary. The measure of one angle is 22 degrees more than the measure of the other angle. Find the measures of each angle.

OpenStudy (anonymous):

is this in a triangle?

OpenStudy (anonymous):

let one angle be x so other = 90-x degrees. x - (90 - x) = 22 2x = 112 x = 56 so the angles are 56 and 56-22 = 34 56 and 34 degrees